OpenStudy (anonymous):

find the x-intercept and y-intercept of the line? 5x+10y=30 i know x-intercept=6 and the y-intercept=3 but i need to know how they got it can some one teach me the steps so i can know how to do it myself please

OpenStudy (igreen):

5x + 10y = 30 First you have to change it to Point Slope Form(y = mx + b) 5x + 10y = 30 Subtract 5x to both sides: 10y = -5x + 30 Divide 10 to both sides of the equation: y = -5/10x + 3 y = -1/2x + 3 So you switched it to Point Slope form. Point Slope Form: y = mx + b Where m = slope And b = y-intercept

OpenStudy (igreen):

So what's the y-intercept? @ObeyBaby_nemo

OpenStudy (anonymous):

3

OpenStudy (igreen):

Yes, that's correct. To find the x-intercept input 0 for 'y'. 5x + 10y = 30 5x + 10(0) = 30 5x = 30 Divide 5 to both sides, and you'll get your x-intercept.

OpenStudy (anonymous):

when yah done igreen could i get some help please

OpenStudy (igreen):

I'm done. What do you need help with? @kfullwood1